OpenStudy (anonymous):

MEDAL FOR THE BEST ANSWER!! What is the value of 10C7? A. 720 B. 120 C. 5040 D. 604,800

OpenStudy (raden):

use the formula : nCr = n!/(r!*(n-r)!)

OpenStudy (anonymous):

\[10C7=10C10-7=10C3=\frac{ 10*9*8 }{3*2*1 }=?\]

OpenStudy (anonymous):

I get thank you for help!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ncr=ncn-r